OpenStudy (prince1342):

2. The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work.

OpenStudy (prince1342):

Ill give a medal

OpenStudy (anonymous):

1. The distance traveled by a falling object is given by the formula d = 0.5gt^2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work. NOTE: Actually, g is the gravitional acceleration - not a force. However, weight, which is caused by gravity is a force. d = 0.5gt^2 d/0.5t^2 = g g = d/0.5t^2 t = 10 sec d = 490 ft g = 9.8 ft/sec² ANSWER NOTE: The magnitude of calculated gravitional acceleration is correct; however, the number shown is g IN METRIC UNITS, whereas, the problem was presented IN U.S. UNITS. Of course, it might be just a coincidence.

OpenStudy (anonymous):

does this help?

OpenStudy (prince1342):

YESSSS YOUR AMAZING